# short-time Fourier transform iSTFT inverse short-time Fourier transform

### v1

@@ Abstract
This paper presents an unsupervised method for single-channel blind dereverberation and room impulse response (RIR) estimation, called BUDDy. The algorithm is rooted in Bayesian posterior sampling: it combines a likelihood model enforcing fidelity to the reverberant measurement, and an anechoic speech prior implemented by an unconditional diffusion model. We design a parametric filter representing the RIR, with exponential decay for each frequency subband. Room acoustics estimation and speech dereverberation are jointly carried out, as the filter parameters are iteratively estimated and the speech utterance refined along the reverse diffusion trajectory. In a blind scenario where the room impulse response is unknown, BUDDy successfully performs speech dereverberation in various acoustic scenarios, significantly outperforming other blind unsupervised baselines. Unlike supervised methods, which often struggle to generalize, BUDDy seamlessly adapts to different acoustic conditions. This paper extends our previous work by offering new experimental results and insights into the algorithm’s performance and versatility. We first investigate the robustness of informed dereverberation methods to RIR estimation errors, to motivate the joint acoustic estimation and dereverberation paradigm. Then, we demonstrate the adaptability of our method to high-resolution singing voice dereverberation, study its performance in RIR estimation, and conduct subjective evaluation experiments to validate the perceptual quality of the results, among other contributions. Audio samples and code can be found online.111uhh.de/sp-inf-buddy
@@
Acoustics, diffusion models, reverberation, speech enhancement.
@@ I Introduction
Reverberation is a natural phenomenon caused by acoustic waves propagating in a room and getting reflected by walls. Reverberation and particularly late reflections often degrade speech intelligibility and quality for normal listeners, and even more severely so for hearing-impaired listeners [1]. Therefore, many communication devices now include a dereverberation algorithm, which aims to recover the anechoic component of speech. This paper considers the scenario in which recordings from only one microphone are available, which is more challenging than a multi-channel scenario [2].
@@
Data-driven approaches rely less on distributional assumptions than statistical methods but instead directly learn the signal properties and structures from data [14]. Among these, supervised predictive models are particularly popular for blind dereverberation: these range from time-frequency masking [15] and mapping[16] to algorithms operating on the cepstrum [17] or directly on the waveform [18, 19].
@@
In contrast, unsupervised methods have been getting little visibility but boast interesting properties such as improved robustness to unseen acoustic conditions without the need for retraining. An unsupervised method for informed single-channel dereverberation based on diffusion models was proposed in our prior work [26]. That approach is based on Bayesian diffusion posterior sampling (DPS) [27], combining a diffusion-based anechoic speech prior and a Gaussian likelihood model for state-of-the-art informed dereverberation. However, as shown in this work, such an informed algorithm is sensitive to even small RIR estimation errors, rendering it impractical in real-life scenarios.
@@
  • In Section II, we investigate the robustness of informed dereverberation approaches in partially blind scenarios. We highlight the limitations of these approaches when the RIR is perturbed with Gaussian noise or estimated blindly using a state-of-the-art RIR estimator [41].
@@
  • Section V-A extends the evaluation of BUDDy for speech dereverberation beyond instrumental metrics, including a subjective listening test and a set of ablation studies.
@@
  • Section V-B presents new experiments on applying BUDDy to singing voice dereverberation at a sampling rate of 44.1 kHz, which is higher compared to the 16-kHz sampling rate used in our speech experiments [40]. The results, which also include a subjective listening test, indicate that our method significantly outperforms existing unsupervised state-of-the-art approaches and performs comparably to a supervised adaptation of the proposed method.
@@
  • Finally, Section V-C assesses BUDDy’s performance in RIR estimation against a state-of-the-art supervised estimator [41]. We use frequency-wise acoustic descriptors to evaluate the accuracy of BUDDy on reverberation time and clarity.
@@
We organize the paper as follows. Section II reports on the robustness of RIR-informed methods, providing context for the proposed approach. In Section III, we introduce diffusion-based generative models and posterior sampling methods for dereverberation using diffusion priors as proposed in previous work [26]. Then in Section IV, we introduce our blind unsupervised dereverberation method BUDDy [40]. The experiments and results mentioned above are presented in Section V. Lastly, Section VI concludes the paper.
II Robustness of RIR-Informed Methods in Partially Blind Scenarios
We consider dereverberation under the prism of inverse problem solving: we wish to retrieve the anechoic utterance waveform $\mathbf{x}_{0}\in\mathbb{R}^{L}$, where $L$ is the speech utterance length, given the reverberant measurement $\mathbf{y}$. Reverberation is often modelled as a convolution between the anechoic speech with a RIR $\mathbf{h}\in\mathbb{R}^{L_{\mathbf{h}}}$, such that $\mathbf{y}=\mathbf{h}\ast\mathbf{x}_{0}$, where $\ast$ is the discrete convolution operator in the time domain, resulting in $\mathbf{y}\in\mathbb{R}^{L+L_{\mathbf{h}}-1}$.
Informed dereverberation algorithms such as [5, 26] assume complete knowledge of the room acoustics as provided by the RIR $\mathbf{h}$. However, as pointed out in Section I, even if the RIR is perfectly known, single channel dereverberation is not trivial as RIRs represent mixed-phase systems such that causal and stable inverse filters do not exist [12]. In practical applications, the RIR is often unknown, and even when it can be estimated, there are typically estimation errors making the task of robust single channel dereverberation even harder. Before delving into blind dereverberation, which is the main focus of this paper, we first examine and demonstrate the sensitivity of informed dereverberation methods in such partially blind scenarios, i.e., when RIRs are known up to estimation errors.
We investigate here two methods, the first of which is our prior work InfDerevDPS [26], which is based on Bayesian diffusion posterior sampling (DPS) [27]. InfDerevDPS combines a diffusion-based anechoic speech prior and a Gaussian likelihood model which measures the adherence of the current estimate to the reverberant utterance, given the reverberation operator. The diffusion-based speech prior and sampling technique for InfDerevDPS are presented in Section III-B hereafter. The second method RIF+Post [5] performs regularized inverse filtering in the Fourier domain, followed by traditional speech enhancement.
We start studying the case where the oracle RIR is corrupted by Gaussian noise. The results displayed in Fig. 1 indicate that the performance of both the diffusion-based and the traditional method dwindles as the noise power increases. This suggests that informed methods have very limited robustness to errors in the provided RIR. This is problematic since perfect sample-wise estimation of RIRs is an arduous problem, given the statistical nature of RIRs [11].
We now shift to a more practical scenario where the RIR is estimated from the reverberant speech only by a DNN-based estimator. In particular, we employ FiNS [41], a state-of-the-art supervised RIR estimator which obtains RIR estimates based on the reverberant utterance (see Section V-C1 for details). We compare in Table I results where the RIR is perfectly known (i.e. informed scenario) versus when it is estimated by FiNS (i.e. partially blind). The acoustic conditions in the considered evaluation set match those of the training set. Therefore, since FiNS was trained in a supervised fashion using paired reverberant/RIR data, it is expected to perform well on such conditions. Indeed, through informal listening, we notice that FiNS produces perceptually reasonable RIR estimates. Yet, the dereverberation performance of both InfDerevDPS and RIF+Post is poor when the RIR is estimated with FiNS [41], as opposed to when the RIR is perfectly known. This shows the limited robustness of these informed methods, and suggests that in such blind case the RIR should be jointly estimated with the anechoic speech. This is the paradigm followed by our method BUDDy, which we will introduce in the next sections.
TABLE I: Dereverberation results on matched reverberant VCTK dataset. We indicate for each method in the table if it operates in a blind scenario.
Method Blind DNS-MOS PESQ
Reverberant - 3.14 $\pm$ 0.52 1.61 $\pm$ 0.37
RIF+Post [5] 3.41 $\pm$ 0.47 2.66 $\pm$ 0.40
InfDerevDPS [26] 3.91 $\pm$ 0.33 3.95 $\pm$ 0.42
FiNS/RIF+Post [41, 5] 2.18 $\pm$ 0.38 1.33 $\pm$ 0.19
FiNS/InfDerevDPS[41, 26] 2.19 $\pm$ 0.43 1.32 $\pm$ 0.18
BUDDy (ours) [40] 3.76 $\pm$ 0.41 2.30 $\pm$ 0.53
III Informed Diffusion-Based Dereverberation
@@
III-A Diffusion-Based Generative Models
@@
Diffusion models [22, 42] have achieved remarkable success across various domains, including speech [43]. They break down the problem of generating high-dimensional complex data into a series of easier denoising tasks. Training a diffusion model first requires defining a forward process, which gradually adds noise to data points, turning the target data distribution into a tractable Gaussian distribution. Conversely, data generation is accomplished by reversing the corruption process. First, an initial sample is drawn from a Gaussian distribution, and then the model iteratively removes noise until a clean sample from the target distribution emerges.
The reverse process, which defines a transport between a Gaussian prior distribution and a target data distribution $p_{\text{data}}$, can be characterized by the probability flow ordinary differential equation (ODE):
$\mathrm{d}\mathbf{x}_{\tau}=[\mathbf{f}(\mathbf{x}_{\tau},\tau)-\frac{1}{2}g^{2}(\tau)\nabla_{\mathbf{x}_{\tau}}\log p(\mathbf{x}_{\tau})]\mathrm{d}\tau\,,$ (1)
@@
where diffusion time $\tau$ flows in reverse from $\tau=T$ to $\tau=T_{\mathrm{min}}\ll T$. The diffusion state $\mathbf{x}_{\tau}\in\mathbb{R}^{L}$ starts from the initial condition $\mathbf{x}_{T}\in\mathbb{R}^{L}$ and ends at $\mathbf{x}_{0}\in\mathbb{R}^{L}\sim p_{\text{data}}$, where $L$ is the length of the time-domain speech utterance. We adopt the parameterization proposed by Karras et al. [44], which defines the drift and diffusion parameters as $f(\mathbf{\mathbf{x}_{\tau}},\tau)=0$ and $g(\tau)=\sqrt{2\tau}$, respectively. Similarly, we adopt $\sigma(\tau)=\tau$ as the noise schedule which defines the so-called transition kernel i.e. the marginal density of the forward process:
$q_{\tau}(\mathbf{x}_{\tau}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{0},\sigma^{2}(\tau)\mathbf{I})\,,$ (2)
@@
where $\mathbf{I}\in\mathbb{R}^{L\times L}$ is the identity matrix. The score function $\nabla_{\mathbf{x}_{\tau}}\log p(\mathbf{x}_{\tau})$ indicates the direction of maximum data likelihood. In practice, it is intractable and we need to estimate it with a score model $\mathbf{s}_{\theta}(\mathbf{x}_{\tau},\tau)$ parameterized with a deep neural network (DNN). Vincent et al. have shown that the score model $\mathbf{s}_{\theta}(\mathbf{x}_{\tau},\tau)$ can be optimized using denoising score matching, i.e. matching the score of the Gaussian transition kernel $q_{\tau}(\mathbf{x}_{\tau}|\mathbf{x}_{0})$ instead of the score of the unknown probability $p(\mathbf{x}_{\tau})$ [45]. The score of the transition kernel $q_{\tau}(\mathbf{x}_{\tau}|\mathbf{x}_{0})$ can be obtained from (2) as:
$\nabla_{\mathbf{x}_{\tau}}q_{\tau}(\mathbf{x}_{\tau}|\mathbf{x}_{0})=\frac{\mathbf{x}_{\tau}-\boldsymbol{\mu}(\mathbf{x}_{0},\tau)}{\sigma(\tau)^{2}}\,.$ (3)
@@
The score model $\mathbf{s}_{\theta}$ is therefore trained using the denoising score-matching objective:[45]

### v2

@@ Abstract
This paper presents an unsupervised method for single-channel blind dereverberation and room impulse response (RIR) estimation, called BUDDy. The algorithm is rooted in Bayesian posterior sampling: it combines a likelihood model enforcing fidelity to the reverberant measurement, and an anechoic speech prior implemented by an unconditional diffusion model. We design a parametric filter representing the RIR, with exponential decay for each frequency subband. Room acoustics estimation and speech dereverberation are jointly carried out, as the filter parameters are iteratively estimated and the speech utterance refined along the reverse diffusion trajectory. In a blind scenario where the RIR is unknown, BUDDy successfully performs speech dereverberation in various acoustic scenarios, significantly outperforming other blind unsupervised baselines. Unlike supervised methods, which often struggle to generalize, BUDDy seamlessly adapts to different acoustic conditions. This paper extends our previous work by offering new experimental results and insights into the algorithm’s versatility. We demonstrate the robustness of our proposed method to new acoustic and speaker conditions, as well as its adaptability to high-resolution singing voice dereverberation, using both instrumental metrics and subjective listening evaluation. We study BUDDy’s performance for RIR estimation and observe it surpasses a state-of-the-art supervised DNN-based estimator on mismatched acoustic conditions. Finally, we investigate the sensitivity of informed dereverberation methods to RIR estimation errors, thereby motivating the joint acoustic estimation and dereverberation design. Audio examples and code can be found online.111uhh.de/sp-inf-buddy
@@
Acoustics, reverberation, speech enhancement.
@@ I Introduction
Reverberation is a natural phenomenon caused by acoustic waves propagating in a space and reflecting off various surfaces, such as walls, ceilings, floors, and objects within the environment. Reverberation and particularly late reflections often degrade speech intelligibility and quality for normal listeners, and even more severely so for hearing-impaired listeners [1]. Therefore, many communication devices now include a dereverberation algorithm, which aims to recover the anechoic component of speech. This paper considers the case in which recordings from only one microphone are available, which is more challenging than a multi-channel case [2].
@@
Data-driven approaches rely less on distributional assumptions than statistical methods but instead directly learn the signal properties and structures from data [14]. Most of these methods are based on supervised learning, where models are trained using paired data. Each input is associated with a corresponding target output, allowing the model to learn a mapping from inputs to outputs. For dereverberation, this typically involves using pairs of anechoic and reverberant speech, where the latter is often produced by convolving anechoic speech signals with RIRs. Supervised predictive models are particularly popular for blind dereverberation: these range from time-frequency masking [15] and mapping[16] to algorithms operating on the cepstrum [17] or directly on the waveform [18, 19].
@@
In contrast, unsupervised methods operate without paired data, relying solely on patterns learned from anechoic speech signals. These approaches have been getting little visibility but boast interesting properties such as improved robustness to unseen acoustic conditions without the need for retraining. An unsupervised method for informed single-channel dereverberation based on diffusion models was proposed in our prior work [26]. That approach is based on Bayesian diffusion posterior sampling (DPS) [27], combining a diffusion-based anechoic speech prior and a Gaussian likelihood model for state-of-the-art informed dereverberation. However, as shown in this work, such an informed algorithm is sensitive to even small RIR estimation errors, rendering it impractical in real-life scenarios.
@@
  • Section IV-A extends the evaluation of BUDDy for speech dereverberation beyond instrumental metrics, including a subjective listening test and a set of ablation studies. Notably, we perform experiments on speech recordings made in real environments, rather than using the convolution model, a scenario in which BUDDy performs especially well.
@@
  • Section IV-B presents new experiments on applying BUDDy to singing voice dereverberation at a sampling rate of 44.1 kHz, which is higher than the 16-kHz sampling rate used in our speech experiments [40]. The results, which also include a subjective listening test, indicate that our method significantly outperforms existing unsupervised state-of-the-art approaches and performs comparably to supervised baselines.
  • In Section IV-C, we investigate the robustness of informed dereverberation approaches in partially blind scenarios, in comparison to BUDDy. We highlight the limitations of these approaches when the RIR is perturbed with Gaussian noise or estimated blindly using a state-of-the-art RIR estimator [41].
@@
  • Finally, Section IV-D assesses BUDDy’s performance in RIR estimation against a state-of-the-art supervised estimator [41]. We use frequency-wise acoustic descriptors to evaluate the accuracy of BUDDy on reverberation time and clarity.
@@
We organize the paper as follows. In Section II, we introduce diffusion-based generative models and posterior sampling methods for informed dereverberation using diffusion priors as proposed in previous work [26]. Then in Section III, we introduce our blind unsupervised dereverberation method BUDDy [40], which extends the posterior sampling method presented in the previous section to the blind scenario where the RIR is not available. The experiments and results mentioned above are presented in Section IV. Section V concludes the paper.
II Informed Diffusion-Based Dereverberation
@@
Throughout this paper, we use the following notations: considering dereverberation under the prism of inverse problem solving, we wish to retrieve the anechoic time-domain utterance $\mathbf{x}_{0}\in\mathbb{R}^{L}$, where $L$ is the length of the utterance, given the reverberant measurement $\mathbf{y}$. As in most dereverberation studies, reverberation is modeled as a convolution between anechoic speech with a RIR $\mathbf{h}\in\mathbb{R}^{L_{\mathbf{h}}}$, such that $\mathbf{y}=\mathbf{h}\ast\mathbf{x}_{0}$, where $\ast$ is the discrete convolution operator in the time domain, resulting in $\mathbf{y}\in\mathbb{R}^{L+L_{\mathbf{h}}-1}$.
II-A Diffusion-Based Generative Models
@@
Diffusion models [22, 42] have achieved remarkable success across various domains, including speech [43]. They break down the problem of generating high-dimensional complex data into a series of easier denoising tasks. Training a diffusion model first requires defining a forward process, which gradually adds noise to data points, turning the target data distribution into a tractable Gaussian distribution.
The forward process is the solution of the following stochastic differential equation (SDE):
$\mathrm{d}\mathbf{x}_{\tau}=\mathbf{f}(\mathbf{x}_{\tau},\tau)\mathrm{d}\tau+g(\tau)\mathrm{d}\mathbf{w}_{\tau}\,,$ (1)
@@
where the diffusion time $\tau$ indexes the stochastic process $\mathbf{x}_{\tau}$ from $\tau=T_{\mathrm{min}}$ to $\tau=T\gg T_{\mathrm{min}}$. The minimal process time $T_{\mathrm{min}}$ is chosen stricly positive to avoid irregularities around 00. The Wiener process $\mathbf{w}_{\tau}$ injects noise with independent and normally distributed increments, that is, $\mathbf{w}_{\tau+\mathop{}\!\mathrm{d}{\tau}}-\mathbf{w}_{\tau}\sim\mathcal{N}(\mathbf{0},\mathop{}\!\mathrm{d}{\tau}\,\mathbf{I})$ where $\mathbf{I}\in\mathbb{R}^{L\times L}$ is the identity matrix [44]. The diffusion state $\mathbf{x}_{\tau}\in\mathbb{R}^{L}$ starts at a clean speech data point $\mathbf{x}_{0}\in\mathbb{R}^{L}\sim p_{\text{data}}$ and ends at the final state $\mathbf{x}_{T}\in\mathbb{R}^{L}$ which contains mostly Gaussian noise. We adopt the parameterization proposed by Karras et al. [45], which defines the drift and diffusion parameters as $f(\mathbf{\mathbf{x}_{\tau}},\tau)=0$ and $g(\tau)=\sqrt{2\tau}$, respectively. This results in a noise schedule $\sigma(\tau)=\tau$ which determines the so-called transition kernel i.e. the marginal density of the forward process [45]
$q_{\tau}(\mathbf{x}_{\tau}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{0},\sigma^{2}(\tau)\mathbf{I})\,.$ (2)
Conversely, data generation is accomplished by reversing the forward corruption process. First, an initial sample is drawn from a Gaussian distribution, and then the model iteratively removes noise until a clean sample from the target distribution emerges. The reverse process, can be characterized by the probability flow ordinary differential equation (ODE), which has the same marginal distributions as the reverse SDE canonically associated to the forward SDE (1) [46]
$\mathrm{d}\mathbf{x}_{\tau}=\left[\mathbf{f}(\mathbf{x}_{\tau},\tau)-\frac{g^{2}(\tau)}{2}\nabla_{\mathbf{x}_{\tau}}\log p(\mathbf{x}_{\tau})\right]\mathrm{d}\tau\,,$ (3)
@@
where diffusion time $\tau$ flows in reverse from $\tau=T$ to $\tau=T_{\mathrm{min}}$. The diffusion state $\mathbf{x}_{\tau}\in\mathbb{R}^{L}$ starts from the initial state $\mathbf{x}_{T}\in\mathbb{R}^{L}$ and ends at $\mathbf{x}_{0}\in\mathbb{R}^{L}\sim p_{\text{data}}$. The score function $\nabla_{\mathbf{x}_{\tau}}\log p(\mathbf{x}_{\tau})$ indicates the direction towards regions of higher probability under the model’s distribution. In practice, it is intractable and we need to estimate it with a score model $\mathbf{s}_{\theta}(\mathbf{x}_{\tau},\tau)$ parameterized with a deep neural network (DNN). Vincent et al. have shown that the score model $\mathbf{s}_{\theta}(\mathbf{x}_{\tau},\tau)$ can be optimized using denoising score matching [47], i.e. matching the score of the Gaussian transition kernel $q_{\tau}(\mathbf{x}_{\tau}|\mathbf{x}_{0})$ instead of the score of the unknown probability $p(\mathbf{x}_{\tau})$. The score of the transition kernel $q_{\tau}(\mathbf{x}_{\tau}|\mathbf{x}_{0})$ can be obtained from (2) as
$\nabla_{\mathbf{x}_{\tau}}\log q_{\tau}(\mathbf{x}_{\tau}|\mathbf{x}_{0})=-\frac{\mathbf{x}_{\tau}-\mathbf{x}_{0}}{\sigma^{2}(\tau)}\,.$ (4)
@@
The score model $\mathbf{s}_{\theta}$ is therefore trained using the denoising score-matching objective [47]
@@
II-B Diffusion Posterior Sampling for Dereverberation
We discuss in this section how diffusion priors can be adapted in order to solve inverse problems. While some traditional methods derive maximum a posteriori estimators for blind dereverberation [9, 10, 11], we exploit the generative nature of diffusion models to solve this inverse problem using posterior sampling. Assuming that the RIR $\mathbf{h}$ is known, we attempt to sample from the posterior distribution of the anechoic speech given the measurement and the RIR $p(\mathbf{x}_{0}|\mathbf{y})\propto p(\mathbf{y}|\mathbf{x}_{0})p(\mathbf{x}_{0})$. Although we do not have an explicit prior $p(\mathbf{x}_{0})$ like in e.g. VAE frameworks, we leverage the implicit prior $p_{\theta^{*}}(\mathbf{x}_{0})$ given by the pretrained diffusion model $\mathbf{s}_{\theta^{*}}(\mathbf{x}_{\tau},\tau)$ where $\theta^{*}$ represents the (fixed) parameters optimizing the training objective (5). Sampling is then achieved by solving the probability flow ODE (3), replacing the unconditional score function by the posterior score $\nabla_{\mathbf{x}_{\tau}}\log p(\mathbf{x}_{\tau}|\mathbf{y})$ [23] obtained through Bayes’ rule
$\mathrm{d}\mathbf{x}_{\tau}$ $=\left[\mathbf{f}(\mathbf{x}_{\tau},\tau)-\frac{1}{2}g^{2}(\tau)\nabla_{\mathbf{x}_{\tau}}\log p(\mathbf{x}_{\tau}|\mathbf{y})\right]\mathrm{d}\tau\,,$
$\approx\left[\mathbf{f}\left(\mathbf{x}_{\tau},\tau\right)-\frac{1}{2}g^{2}(\tau)\left(\mathbf{s}_{\theta^{*}}(\mathbf{x}_{\tau},\tau)+\nabla_{\mathbf{x}_{\tau}}\log p(\mathbf{y}|\mathbf{x}_{\tau})\right)\right]\mathrm{d}\tau\,.$ (6)
The likelihood score $\nabla_{\mathbf{x}_{\tau}}\log p(\mathbf{y}|\mathbf{x}_{\tau})$ is in general intractable for $\tau>0$. Following [27], we employ a plug-in estimate of $\mathbf{x}_{0}$ denoted as $\hat{\mathbf{x}}_{0}^{\tau}$ which we derive using Tweedie’s formula, i.e. one-step denoising of $\mathbf{x}_{\tau}$ using the diffusion model
$\hat{\mathbf{x}}_{0}^{\tau}\overset{\Delta}{=}\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{\tau}]\approx\mathbf{x}_{\tau}+\sigma^{2}(\tau)\mathbf{s}_{\theta^{*}}(\mathbf{x}_{\tau},\tau)\,.$ (7)
We assume that this estimate is a sufficient statistic for $\mathbf{x}_{\tau}$, which results in a first assumption $p(\mathbf{y}|\mathbf{x}_{\tau})\approx p(\mathbf{y}|\hat{\mathbf{x}}_{0}^{\tau})$.
In order to approximate $p(\mathbf{y}|\hat{\mathbf{x}}_{0}^{\tau})$, previous work [26] models the error between $\mathbf{y}$ and its estimation to follow a zero-mean Gaussian distribution in the time domain. The corresponding expression for the likelihood score $\nabla_{\mathbf{x}_{\tau}}\log p(\mathbf{y}|\mathbf{x}_{\tau})$ is then a simple weighted $L^{2}$-distance between $\mathbf{y}$ and $\mathbf{h}\ast\hat{\mathbf{x}}_{0}^{\tau}$. However far better dereverberation performance and speech quality can be achieved by substituting the obtained distance with a $L^{2}$-distance between compressed short-time Fourier transform (STFT) representations instead. This is analogous to modeling the likelihood as
$p(\mathbf{y}|\mathbf{x}_{0})\propto\mathrm{exp}\big{(}-\zeta(\tau)\,\mathcal{C}(\mathbf{y},\mathbf{h}\ast\hat{\mathbf{x}}_{0}^{\tau})\,\big{)}\,,$ (8)
where the scalar $\zeta(\tau)$ controls the influence of the likelihood score term during sampling, and $\mathcal{C}(\cdot,\cdot)$ is the cost function
$\mathcal{C}(\mathbf{u},\mathbf{v})=\frac{1}{M}\sum_{m=1}^{M}\sum_{k=1}^{K}\rVert S_{\mathrm{comp}}(\mathbf{u})_{m,k}-S_{\mathrm{comp}}(\mathbf{v})_{m,k}\lVert_{2}^{2}\,.$ (9)
There, $S_{\mathrm{comp}}(\mathbf{u})\in\mathbb{C}^{M\times K}$ denotes the magnitude-compressed STFT of $\mathbf{u}$, comprising $M$ time frames and $K$ frequency bins
$S_{\mathrm{comp}}(\mathbf{u})=|\text{STFT}(\mathbf{u})|^{2/3}\exp{j\angle\text{STFT}(\mathbf{u})}\,.$ ) . (10)
We apply this compression to boost low-energy components as typically observed in high frequencies of speech signals or in late reverberation tails, and account for the heavy-tailedness of speech distributions [4]. Such a strategy is also employed in [48] for data representation. In our case, we only use this non-linear transformation in the cost function (9), whereas the diffusion process itself uses the original time-domain representation.
$\zeta(\tau)=\frac{\sqrt{L}\,\tilde{\zeta}}{\sigma(\tau)\norm{\nabla_{\mathbf{x}_{\tau}}\mathcal{C}\left(\mathbf{y},\mathbf{h}\ast\hat{\mathbf{x}}_{0}(\mathbf{x}_{\tau})\right)}_{2}}\,,$ (11)
where $\tilde{\zeta}$ is a fixed coefficient.
The resulting informed dereverberation algorithm is a slight variation of our previous work [26]. In the following we refer to this approach as InfDerevDPS.
III Blind Diffusion-Based Dereverberation